# Enhanced detection of time-dependent dielectric structure: Rayleigh limit and quantum vacuum

### v3

@@ Abstract
[…]
@@ I Introduction
Inverse optics determines the dielectric susceptibility of inhomogeneous materials by shining (incident) light with known characteristics on them and detecting scattered light [1, 2]. Recent advances in inverse optics relate to the use of quantum features of incident light [3, 4, 5, 6, 7, 8]. In particular, quantum features improve on the classical Rayleigh limit that bounds the far-field resolution of dielectric susceptibility, given the wavelength of the incident light. Improvement is possible (up to eight times for two-photon light [8]), but it connects with difficulties of preparing specific quantum states of light and with having prior information on the dielectric sample [8].
We are looking for additional resources that will allow for better resolution. Our results identify one such resource that relies on the dielectric’s motion or on imitating this motion via time modulation of an inhomogeneous dielectric. Our proposal is dual to Doppler metrology [9]: we set the moving object in motion to analize its internal structure, rather than finding the velocity of the object.
@@
Electrodynamics of relativistically moving bodies has traditionally focused on electron clouds or plasma jets [20, 21, 14]. A recent activity in creating metamaterials with a space-time-modulated electric susceptibility renewed interest in this field [22, 23, 24]. Though these metamaterials do not move, they are expected to mimic moving dielectrics.
Here we aim to show that motion or time modulation of an inhomogeneous dielectric is a resource for improving resolution of its internal structure, i.e. seeing deeper into the dielectric structure. It turns out that for a modulated or moving dielectric, incident light is not even necessary: the object reveals its structure through the vacuum response of photodetector. This effect comes from negative frequencies of the quantum electric field. More specifically, we develop scattering theory for two situations:
(i) A spatially and temporally modulated metamaterial, whose dielectric susceptibility is similar to that of moving dielectrics.
(ii) A dielectric moving with a constant velocity, a problem we studied within relativistic optics.
For (i) this vacuum response takes place in the far-field domain of spherical waves. In its specific realizations (broadband harmonic modulation) it has certain analogies to the anomalous Doppler effect; see [25] for a review on this effect. For (ii) the vacuum photodetection is limited to a near-field zone of cylindrical waves, but it can be a long-range effect if the motion is ultra-relativistic. When shining is present, we show that (i) can overcome the classical Rayleigh resolution limit in the far-field situation. This allows for an enhanced detection of dielectric features along the direction of modulation. Technically, this means that the Fourier image of the dielectric susceptibility is determined at wave vectors sizably larger than $\omega/c$, where $\omega$ is the photodetection frequency.
This paper is organized as follows. Sections II and III study quantum scattering theory for a space-time modulated dielectric. Both vacuum response (section III.1) and one-photon shining (section III.3) are considered. Section IV developed a relativistic scattering theory for a moving dielectric, accounting for both spatial and temporal dispersion (though we eventually focus on temporal dispersion only). In contrast to the phenomenological approach of sections II and III, the relativisitic consideration is more based on first principles. The last section summarizes.
@@ II Scattering theory for inhomogeneous dielectrics
We assume that the electromagnetic field is quantized making (2) the Heisenberg equation for the operator $E_{\beta}$ [26]. But the matter is classical and has the average dielectric susceptibility $\varepsilon\left(\mathbf{r,}t\right)-1$ [26]. Hence, in (1) we left aside both the time-dispersion of dielectrics and its quantum features. Both are straightforward to include: time-dispersion will amount to an integral equation in (1), while the quantum features can be introduced via making $\varepsilon\left(\mathbf{r,}t\right)$ an operator and postponing the averaging over the dielectric state till the final results [26]. We avoid introducing these complications because in the present phenomenological set-up, none of them are essential to our main results. For the same reason, we neglected anisotropy in $\varepsilon\left(\mathbf{r,}t\right)$. In section IV, we will study a relativistically moving dielectric, and we will need to account (at least) for time-dispersion features.
@@
where $\mathbf{e}_{\lambda}\left(\mathbf{q}\right)=\mathbf{e}_{\lambda}\left(-\mathbf{q}\right)$ is the real polarisation vector, $A+{\rm h.c.}=A+A^{\dagger}$, $\delta_{\lambda\lambda^{\prime}}$ is Kroenecker’s delta, and ${a}_{\lambda}(\mathbf{q})$ is the annihilation operator. Note that ${a}_{\lambda}(\mathbf{q})$ and ${a}_{\lambda^{\prime}}(\mathbf{q}^{\prime})$ commute. Eqs. (9, 10) define orthogonality features of polarization vectors, where repeated $\lambda$ implies summation. Now $\mathbf{{E}}^{[\text{in}]}(\mathbf{r},t)$ contains contributions both from creation ${a}^{\dagger}_{\lambda}\left(\mathbf{q}\right)$ and annihilation ${a}_{\lambda}\left(\mathbf{q}\right)$ operators at (resp.) positive and negative frequencies, while only ${a}_{\lambda}\left(\mathbf{q}\right)$ contribute to $\mathbf{{E}}^{[\text{in}]}[\omega>0,\mathbf{r}]$.
@@
$G_{\alpha\beta}[\omega>0,\mathbf{r}]=\Big{(}\delta_{\alpha\beta}+\frac{\partial_{\alpha}\partial_{\beta}}{\omega^{2}}\Big{)}\frac{e^{i\omega r}}{(-4\pi)r},~{}r=|\mathbf{r}|.$ (16)
@@
Note that for the time-independent dielectric, $\bar{\varepsilon}(\omega)\propto\delta(\omega)$, and only the annihilation operator containing part survives in (17, 5), because (5) refers to $\omega>0$, and then $V_{\alpha\beta}[\mathbf{r},-\mathbf{q},\omega+q]=0$ drops out. More generally, also the creation operator will contribute to (5), which is the core of our effects. We now apply (17, 5) to a resting, but space-time-modulated dielectric.
@@ III Space-time-modulated dielectric
Figure 1: Scattering from resting, space-time-modulated dielectrics (modulation devices are not shown). Shadowed domain denotes the incident field $\mathbf{E}^{[{\rm in}]}$. This can be a single-photon field with momentum $\mathbf{k}$ or the vacuum state of the field; cf. (34). $|\psi\rangle$ is the state of the field, which appears in the photodetection result (5). $\mathbf{n}$ is the unit vector towards the photodetector D that is placed in the far-field domain; see (23). In that domain, only the spherical scattered field $\mathbf{E}^{[{\rm s}]}$ is present if $|\mathbf{n\cdot k}|$ is sufficiently far from $|\mathbf{k}|$, i.e. no backscattering or forward scattering is detected.
@@ III.1 Vacuum response
$\bar{\varepsilon}[\omega,\mathbf{r}]=\chi(\mathbf{r})\left(\delta\left(\omega\right)+\frac{1}{u}\eta[-\omega/u]e^{i\omega x/u}\right),$ (19)
where $\chi\left(\mathbf{r}\right)$ is a localized function that defines the overall shape of the dielectric, $\eta\left(x-ut\right)$ refers to the unknown structure along the $x$-axis that defines the modulation, and $\eta[\omega]$ is the Fourier transform of $\eta(t)$ [cf. (6)]. The modulation speed $u>0$ can be larger than $c=1$ since it does not refer to the energy or information transfer [23, 24]. However, we shall not need $u>c=1$.
We calculate the vacuum response of the photodetector starting from (5, 17, 10) for $|\psi\rangle=|0\rangle$:
@@
If $\chi(\mathbf{r}^{\prime})$ is well-localized e.g. around $\mathbf{r}^{\prime}\simeq 0$, (22) can be simplified further assuming that the photodetector in (21) is placed far away from the dielectric. In this far-field domain $|\mathbf{r}|\gg|\mathbf{r}^{\prime}|$, the photodetector sees a spherical wave [see Appendix A and recall $r=|\mathbf{r}|$],
@@ III.2 Example
×$~{}\,\quad\qquad\times\Big{(}1+n_{x}^{2}y^{2}+\frac{(1-y^{2})(1-n_{x}^{2})}{2}\Big{)}$ (32)
@@ III.2.1 Relations with the anomalous Doppler effect
When looking at (21, 30) it is seen that a non-zero result is achieved only when the modulation speed $u$ is larger than $\omega a$, where $\omega$ is the frequency of the detected photon. The difference $\frac{u}{a}-\omega$ is the frequency $cq=c|\mathbf{q}|$ over which the integration in (21) goes. Recall that this integration came from negative frequencies in (20). This suggests a relation with the anomalous Doppler effect, a mechanism by which an (internally) equilibrium body moving in a medium with a speed larger than the phase velocity of light in that medium can radiate [25]. Cherenkov radiation is a limiting case of the anomalous Doppler effect [13, 20].
@@ III.3 One-photon incident field
We return to (1719) and shine the modulated dielectric with a single-photon state
@@
where $\left|0\right\rangle$ is the vacuum state of the field, and the second relation in (34) ensures $\left\langle\psi|\psi\right\rangle=1$. While the single-photon state was taken for clarity, similar results are obtained for a coherent state of the field that models laser light. Now the photodetector will be placed in the far-field domain; see Fig. 1. Hence, in (5) we can take $\mathbf{E}=\mathbf{E}^{[\text{s}]}$, i.e. only the scattered field contributes. Using (17, 5) we find
@@
where the vacuum contribution (21, 22) enters additively, and where the non-vacuum contribution $\xi_{\alpha}\xi^{*}_{\alpha}$ comes from the positive frequencies. We now assume that the single-photon state (34) has a well-defined momentum $\mathbf{k}$, which factorizes from the polarization:
@@
To calculate (35) in the far-field domain we use (23) and confirm that generally, all frequencies appear in (35). This contrasts with the weak scattering of a resting non-modulated dielectric, where $\mathbf{E}^{[{\rm s}]}$ contains only the incident frequency $|\mathbf{k}|$, hence the photodetector should be tuned to that frequency: $\omega\simeq|\mathbf{k}|$. This fact is seen from (19) after dropping out the modulation contribution. The classical Rayleigh limit relates to this [8].
For clarity, we assume $|\mathbf{k}|\not=\omega$ in (35), i.e. the detection frequency differs from the incident frequency. Together with (36), this allows us to exclude the term $\propto\delta(\omega)$ in (19). We find from (35, 36, 37, 17):
@@ IV.1 Scattering within relativistic electrodynamics
Consider a (non-magnetic) dielectric which in the laboratory frame moves along $x$-axis with a constant velocity $v$; see Fig. 2. We study this problem within the framework of relativistic electrodynamics [11, 29], which will allow a consistent treatment of space-time-dispersion. Relativistic consideration was not needed for the modulated situation (18, 19), because the modulation speed (not limited by special relativity, since there is no energy transfer or information transfer) can be comparable to or even greater than the speed of light [23, 24].
We set $g_{ik}={\rm diag}[1,-1,-1,-1]$ and $c=\hbar=1$. Latin indices assume values $0,1,2,3$ and refer to 4-vectors [cf. (4)]:
@@
where $A^{i}$ is the 4-vector-potential, satisfies the following wave-equation [11, 12] (see Appendix D)
$-\partial_{i}F^{ik}(x)=\partial_{i}\int{\rm d}^{4}\hat{x}\,\bar{\varepsilon}(x,\hat{x})\,(F^{is}u_{s}u^{k}-F^{ks}\,u_{s}\,u^{i})(\hat{x}),$
@@
where $u^{i}$ is the 4-vector of velocity [29], and $\bar{\varepsilon}=\varepsilon-1$ is the dielectric susceptibility.
Eq. (43) is written in the laboratory frame, where the dielectric moves with velocity $v$. To determine the form of $\bar{\varepsilon}(x,\hat{x})$, which contains both inhomogeneity in space, and space-time dispersion, we need to look at the rest frame of the dielectric with coordinates $x^{\prime i}$:
$\bar{\varepsilon}[x^{\prime i}-\hat{x}^{\prime i};x^{\prime i}],$ (44)
where (45) refers to causality, and where the laboratory frame $\bar{\varepsilon}(x,\hat{x})$ in (IV.1) is obtained from (44) after making in (44) the Lorentz transformation
@@
where $q^{l}=(|\mathbf{q}|,\mathbf{q})$, $q_{l}q^{l}=0$, and we where denoted
@@
$A^{{\rm[s]}}_{y}[\omega,\mathbf{r}]=\frac{\gamma}{2iv\pi}\int{\rm d}\mathbf{q}\,C_{y\,\lambda}(\mathbf{q})\times$
$\{a_{\lambda}(\mathbf{q})\,{\cal V}[\mathbf{q},\gamma(q-vq_{x}),\omega-q,\mathbf{r}]$
@@
${\cal V}[\mathbf{q},\gamma(q-vq_{x}),\omega-q,\mathbf{r}]=\int{\rm d}\mathbf{r}^{\prime}\,G[\omega,\mathbf{r}-\mathbf{r}^{\prime}]\times$
$e^{i\mathbf{q}\mathbf{r}^{\prime}+\frac{i(\omega-q)x^{\prime}}{v}}\,\,\widetilde{\varepsilon}[\gamma(q-vq_{x}),-\frac{\omega-q}{\gamma v},y^{\prime},z^{\prime}],$ (55)
where $\widetilde{\varepsilon}$ is the Fourier-transform of (53):
$\widetilde{\varepsilon}[\gamma(q-vq_{x}),-\frac{\omega-q}{\gamma v},y,z]=\int\frac{{\rm d}x}{2\pi}e^{-\frac{ix(\omega-q)}{\gamma v}}\times$
$\hat{\varepsilon}[\gamma(q-vq_{x}),x,y,z],$ (56)
$G[\omega>0,\mathbf{r}]=\frac{e^{i\omega r}}{(-4\pi)r},~{}r=|\mathbf{r}|,$ (57)
and where $G[\omega>0,\mathbf{r}]$ is the scalar Green function; cf. (16). Structurally, (54) is similar to (17). Several differences relate to polarization factors (plus the difference in Green’s functions), but the major difference will be in the far-field radiation, which is absent in (54); see below.

### v4

@@ Abstract
Detection of scattered light can determine the susceptibility of dielectrics. Such imaging normally holds Rayleigh’s limit: details finer than the wavelength of the incident light cannot be determined from the far-field zone. We show that time-modulation of an inhomogeneous dielectric can be used to determine its susceptibility. To this end, we focus on the inverse quantum optics problem for spatially and temporally modulated metamaterial, whose dielectric susceptibility is similar to moving dielectrics. We show that the vacuum contribution to the photodetection signal is non-zero due to the negative frequencies even in the far-field zone. Hence, certain dielectric features can be determined without radiating any incident field on the dielectric. When the incident light is scattered (or re-radiated), the determination of dielectric susceptibility is enhanced and overcomes Rayleigh’s limit in the far-field zone. We study similar effects for an inhomogeneous dielectric moving with a constant speed, a problem we consider within relativistic optics. Now the vacuum contribution to the photodetection signal reflects dielectric features, can be long-range in space but is not a far-field effect.
@@ I Introduction
Inverse optics determines the dielectric susceptibility of inhomogeneous materials by irradiating them with (incident) light with known characteristics and detecting scattered light [1, 2]. Recent advances in inverse optics relate to the use of quantum features of incident light [3, 4, 5, 6, 7, 8]. In particular, quantum features improve on the classical Rayleigh limit that bounds the far-field resolution of dielectric susceptibility, given the wavelength of the incident light. Improvement is possible (up to eight times for two-photon light [8]), but it connects with difficulties of preparing specific quantum states of light and with having prior information on the dielectric sample [8].
We are looking for additional resources that will allow for better resolution. Our results identify a resource that relies on space-time modulation of an inhomogeneous dielectric. As a result of the modulation, the dielectric has a space-time-dependent susceptibility similar to that of a moving dielectric. Our proposal is dual to Doppler metrology [9]: we set the moving object in motion to analyze its internal structure, rather than to find the velocity of the object.
@@
Electrodynamics of relativistically moving bodies has traditionally focused on electron clouds or plasma jets [20, 21, 14]. A recent activity in creating metamaterials with a space-time-modulated electric susceptibility renewed interest in this field [22, 23, 24]. Though these metamaterials do not move, they to a certain extent mimic moving dielectrics.
Here we aim to show that motion or time modulation of an inhomogeneous dielectric is a resource for improving the resolution of its internal structure, i.e. seeing deeper into the dielectric structure. It turns out that for a modulated or moving dielectric, incident light is not even necessary: the object reveals its structure through the vacuum response of the photodetector. This effect comes from up-conversion of field modes/operators from negative to positive frequencies. Specifically, we show that the vacuum response of the far-field photodetector can be used to determine the characteristic length of the dielectric susceptibility for moving dielectrics.
In section III, we develop scattering theory for a spatially and temporally modulated metamaterial, whose dielectric susceptibility is similar to that of moving dielectrics. We show that in this case there is a vacuum response (the one taking place without any incident light), and that this response can take place in the far-field zone of spherical waves. In its specific realizations (broadband harmonic modulation) it has certain analogies to the anomalous Doppler effect; see [25, 26] for reviews on this effect. When incident light and its scattering are present, we show that far-field detection can overcome the classical Rayleigh resolution limit. This allows for an enhanced detection of dielectric features along the direction of modulation. Technically, this means that the Fourier image of the dielectric susceptibility is determined via the far-field photodetection (which measures the intensity of the quantum electric field) at wave vectors sizably larger in modulus than $\omega/c$, where $\omega$ is the photodetection frequency.
The above results concern the space-time modulated dielectric, a behavior that mimics rectilinear motion. To understand precise relations between these two situations, in section IV we study – within relativistic optics – a dielectric moving with a constant velocity. Here the vacuum photodetection is limited to a near-field zone of cylindrical waves, but it can be a long-range effect if the motion is ultra-relativistic. Long-range means that the signal intensity decays as a power law (not exponent) with the distance from the trajectory of the moving dielectric. When cylindrical waves are in the far-field, their intensity decays as the inverse of the distance.
This paper is organized as follows. Sections II and III study quantum scattering theory for a space-time modulated dielectric. The vacuum response is studied in section III.1. In section III.3 we research the modulated situation under one-photon irradiation. Section IV develops a relativistic scattering theory for a moving dielectric, accounting for both spatial and temporal dispersion (though we eventually focus on temporal dispersion only). In contrast to the phenomenological approach of sections II and III, the relativistic consideration of section IV is based on first principles. Section IV can be read independently from sections II and III. The last section summarizes.
@@ II Scattering theory for inhomogeneous dielectrics
We assume that the electromagnetic field is quantized making (2) the Heisenberg equation for the operator $E_{\beta}$ [27]. But the matter is classical and has the average dielectric susceptibility $\varepsilon\left(\mathbf{r,}t\right)-1$ [27]. Hence, in (1) we left aside both the time-dispersion of dielectrics and its quantum features. Both are straightforward to include: time-dispersion will amount to an integral equation in (1), while the quantum features can be introduced by promoting $\varepsilon\left(\mathbf{r,}t\right)$ tos an operator and postponing the averaging over the dielectric state till the final results [27]. We avoid introducing these complications because in the present phenomenological set-up, none of them are essential to our main results. For the same reason, we neglect anisotropy in $\varepsilon\left(\mathbf{r,}t\right)$. In section IV, we study a relativistically moving dielectric, and we need to account (at least) for time-dispersion features.
@@
where $\mathbf{e}_{\lambda}\left(\mathbf{q}\right)=\mathbf{e}_{\lambda}\left(-\mathbf{q}\right)$ is the real polarisation vector, $A+{\rm h.c.}=A+A^{\dagger}$, $\delta_{\lambda\lambda^{\prime}}$ is the Kroenecker delta, and ${a}_{\lambda}(\mathbf{q})$ is the annihilation operator. Note that ${a}_{\lambda}(\mathbf{q})$ and ${a}_{\lambda^{\prime}}(\mathbf{q}^{\prime})$ commute. Eqs. (9, 10) define orthogonality features of polarization vectors, where repeated $\lambda$ implies summation. Now $\mathbf{{E}}^{[\text{in}]}(\mathbf{r},t)$ contains contributions both from creation ${a}^{\dagger}_{\lambda}\left(\mathbf{q}\right)$ and annihilation ${a}_{\lambda}\left(\mathbf{q}\right)$ operators at (resp.) positive and negative frequencies, while only ${a}_{\lambda}\left(\mathbf{q}\right)$ contribute to $\mathbf{{E}}^{[\text{in}]}[\omega>0,\mathbf{r}]$.
@@
$G_{\alpha\beta}[\omega>0,\mathbf{r}]=\Big{(}\delta_{\alpha\beta}+\frac{\partial_{\alpha}\partial_{\beta}}{\omega^{2}}\Big{)}G[\omega,\mathbf{r}],~{}r=|\mathbf{r}|,$ (16)
$G[\omega,\mathbf{r}]=\frac{e^{i\omega r}}{(-4\pi)r},$ (17)
where $G[\omega,\mathbf{r}]$ in (16) is the scalar Green’s function.
@@
Note that for the time-independent dielectric, $\bar{\varepsilon}(\omega)\propto\delta(\omega)$, and only the annihilation operator containing part survives in (18, 5), because (5) refers to $\omega>0$, and then $V_{\alpha\beta}[\mathbf{r},-\mathbf{q},\omega+q]=0$ drops out. More generally, also the creation operator will contribute to (5), which is the core of our effects. We now apply (18, 5) to a space-time-modulated dielectric at rest.
@@ III Space-time-modulated dielectric
Figure 1: Scattering from resting, space-time-modulated dielectrics (modulation devices are not shown). Shadowed domain denotes the incident field $\mathbf{E}^{[{\rm in}]}$. This can be a single-photon field with momentum $\mathbf{k}$ or the vacuum state of the field; cf. (35). $|\psi\rangle$ is the state of the field, which appears in the photodetection result (5). $\mathbf{n}$ is the unit vector towards the photodetector D that is placed in the far-field zone; see (24). In that zone, only the spherical scattered field $\mathbf{E}^{[{\rm s}]}$ is present if $|\mathbf{n\cdot k}|$ is sufficiently far from $|\mathbf{k}|$, i.e. no backscattering or forward scattering is detected.
@@ III.1 Vacuum response
$\bar{\varepsilon}[\omega,\mathbf{r}]=\chi(\mathbf{r})\left(\delta\left(\omega\right)+\frac{1}{u}\eta[\omega/u]e^{i\omega x/u}\right),$ (20)
where $\chi\left(\mathbf{r}\right)$ is a localized function that defines the overall shape of the dielectric, $\eta\left(x-ut\right)$ refers to the unknown structure along the $x$-axis that defines the modulation, and $\eta[\omega/u]$ is the Fourier transform of $\eta(x-ut)$ at the point $\omega/u$ [cf. (6)]. The modulation speed $u>0$ can be larger than $c=1$ since it does not refer to the energy or information transfer [23, 24]. However, we shall not need $u>c=1$.
We calculate the vacuum response of the photodetector starting from (5, 18, 10) for $|\psi\rangle=|0\rangle$:
@@
If $\chi(\mathbf{r}^{\prime})$ is well-localized e.g. around $\mathbf{r}^{\prime}\simeq 0$, (23) can be simplified further assuming that the photodetector in (22) is placed far away from the dielectric. In this far-field zone $|\mathbf{r}|\gg|\mathbf{r}^{\prime}|$, the photodetector sees a spherical wave [see Appendix A and recall $r=|\mathbf{r}|$],
@@ III.2 Example
×$~{}\,\quad\qquad\quad\times\Big{(}1+n_{x}^{2}y^{2}+\frac{(1-y^{2})(1-n_{x}^{2})}{2}\Big{)}$ (33)
@@ III.2.1 Relations with the anomalous Doppler effect
When looking at (22, 31) it is seen that a non-zero result is achieved only when the modulation speed $u$ is larger than $\omega a$, where $\omega$ is the frequency of the detected photon. The difference $\frac{u}{a}-\omega$ is the frequency $cq=c|\mathbf{q}|$ over which the integration in (22) goes. Recall that this integration came from negative frequencies in (21). This suggests a relation with the anomalous Doppler effect, a mechanism by which an (internally) equilibrium body moving in a medium with a speed larger than the phase velocity of light in that medium can radiate [25, 26]. Cherenkov radiation is a limiting case of the anomalous Doppler effect [13, 20].
@@ III.3 One-photon incident field
We return to (1820) and radiate the modulated dielectric with a single-photon state
@@
where $\left|0\right\rangle$ is the vacuum state of the field, and the second relation in (35) ensures $\left\langle\psi|\psi\right\rangle=1$. While the single-photon state was taken for clarity, similar results are obtained for a coherent state of the field that models laser light. As the photodetector is placed in the far-field zone (see Fig. 1), in (5) we can take $\mathbf{E}=\mathbf{E}^{[\text{s}]}$, i.e. only the scattered field contributes. Using (18, 5) we find
@@
where the vacuum contribution (22, 23) enters additively, and where the non-vacuum contribution $\xi_{\alpha}\xi^{*}_{\alpha}$ comes from the positive frequencies. We now assume that the single-photon state (35) has a well-defined momentum $\mathbf{k}$, which factorizes from the polarization:
@@
Eq. (36) confirms that, generally, all frequencies appear in $I[\omega,\mathbf{r};|\psi\rangle\langle\psi|]$. This contrasts with the weak scattering of a non-modulated dielectric at rest, where $\mathbf{E}^{[{\rm s}]}$ contains only the incident frequency $|\mathbf{k}|$, hence the photodetector should be tuned to that frequency: $\omega\simeq|\mathbf{k}|$. This fact is seen from (20) after dropping out the modulation contribution. This fact is related to the classical Rayleigh limit [8].
For clarity, we assume $|\mathbf{k}|\not=\omega$ in (36), i.e. the detection frequency differs from the incident frequency. Together with (37), this allows us to exclude the term $\propto\delta(\omega)$ in (20). We find from (36, 37, 38, 18):
@@ IV.1 Scattering within relativistic electrodynamics
Consider a (non-magnetic) dielectric which in the laboratory frame moves along $x$-axis with a constant velocity $v$; see Fig. 2. We study this problem within the framework of relativistic electrodynamics [11, 30], which will allow a consistent treatment of space-time-dispersion. Relativistic consideration was not needed for the modulated situation (19, 20), because the modulation speed (not limited by special relativity, since there is no energy transfer or information transfer) can be comparable to or even greater than the speed of light [23, 24].
We set the following standard notations:
@@
where $A^{i}$ is the 4-vector-potential, satisfies the following wave-equation [11, 12] (see Appendix D, Eqs. (122, 125))
$-\partial_{i}F^{ik}(x)$ $=\partial_{i}\int{\rm d}^{4}\hat{x}\,\bar{\varepsilon}(x,\hat{x})$
×$\times(F^{is}u_{s}u^{k}-F^{ks}\,u_{s}\,u^{i})(\hat{x}),$ (44)
@@
where $u^{i}$ is the 4-vector of velocity of the dielectric [30], and $\bar{\varepsilon}=\varepsilon-1$ is the dielectric susceptibility. Eqs. (122, 125) were deduced for a dielectric without space-time dispersion, but its generalization towards (44) is straightforward, as it respects the relativistic covariance. Note that $\bar{\varepsilon}(x,\hat{x})$ in (44) contains both inhomogeneity in space, and space-time dispersion.
@@
Now we need to determine the form of $\bar{\varepsilon}(x,\hat{x})$ in (46). Recall that this equation is written in the laboratory frame, where the dielectric moves with velocity $v$. We need to look at the rest frame of the dielectric with coordinates $x^{\prime i}$. Our first assumption is that in the rest frame the space inhomogeneity separates from the space-time dispersion:
$\bar{\varepsilon}(x^{\prime},\hat{x}^{\prime})=\bar{\varepsilon}[x^{\prime i}-\hat{x}^{\prime i};x^{\prime i}],$ (47)
where (48) refers to causality. Next, we assume that the inhomogeneity in (47) refers to the space only, while the dispersion refers only to the time (i.e. the proper space-dispersion is absent): $\bar{\varepsilon}[x^{\prime i}-\hat{x}^{\prime i};x^{\prime i}]=\bar{\varepsilon}[t^{\prime}-\hat{t}^{\prime};\mathbf{r}^{\prime}]$. Now in the laboratory frame $\bar{\varepsilon}(x,\hat{x})$ in (44) is obtained from (47) after making in (47) the Lorentz transformation
@@
where $q^{l}=(|\mathbf{q}|,\mathbf{q})$, $q_{l}q^{l}=0$, and where we denoted
@@
$A^{{\rm[s]}}_{y}[\omega,\mathbf{r}]=\frac{\gamma}{2iv\pi}\int{\rm d}\mathbf{q}\,C_{y\,\lambda}(\mathbf{q})$
×$\times\{a_{\lambda}(\mathbf{q})\,{\cal V}[\mathbf{q},\gamma(q-vq_{x}),\omega-q,\mathbf{r}]$
@@
${\cal V}[\mathbf{q},\gamma(q-vq_{x}),\omega-q,\mathbf{r}]=\int{\rm d}\mathbf{r}^{\prime}\,G[\omega,\mathbf{r}-\mathbf{r}^{\prime}]$
×$\times e^{i\mathbf{q}\mathbf{r}^{\prime}+\frac{i(\omega-q)x^{\prime}}{v}}\,\,\widetilde{\varepsilon}[\gamma(q-vq_{x}),-\frac{\omega-q}{\gamma v},y^{\prime},z^{\prime}],$ (57)
where $G[\omega,\mathbf{r}-\mathbf{r}^{\prime}]$ is defined in (17), and where $\widetilde{\varepsilon}$ is the Fourier-transform of (55):
$\widetilde{\varepsilon}[\gamma(q-vq_{x}),-\frac{\omega-q}{\gamma v},y,z]$ $=\int\frac{{\rm d}x}{2\pi}e^{-\frac{ix(\omega-q)}{\gamma v}}$
×$\times\hat{\varepsilon}[\gamma(q-vq_{x}),x,y,z].$ (58)
Structurally, (56) is similar to (18). Several differences relate to polarization factors (plus the difference in Green’s functions), but the major difference will be in the far-field radiation, which is absent in (56); see below.